OpenStudy (anonymous):

The measure of angle A is twice that of angle B and one less than that of angle C. If the sum of the measures of the three angles is at least 66, what is the least possible measure of angle A? need the solution please. :)

OpenStudy (bahrom7893):

A = 2B A = C-1 2B = C - 1 C-2B = 1 A + B + C is at least 66.. hmm

OpenStudy (anonymous):

then? :)

OpenStudy (anonymous):

\[A+\frac{1}{2}A +A+1<66\]

OpenStudy (anonymous):

solve for A via \[\frac{5}{2}A+1<66\] \[\frac{5}{2}A<65\] \[A<\frac{2}{5}\times 65\] \[A<26\]

OpenStudy (bahrom7893):

sorry i couldnt find this again..

OpenStudy (anonymous):

steps clear?

OpenStudy (anonymous):

where did you get the 5 in 5/2?

OpenStudy (anonymous):

@satellite. where did the 5 came from?

OpenStudy (anonymous):

C=1-2x?

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

i want to write everything in terms of A \[A=2B\] so \[B=\frac{A}{2}\] \[A=C-1\] so \[A+1=C\]

OpenStudy (anonymous):

add them up to get \[A+\frac{1}{2}A+A+1<66\]

OpenStudy (anonymous):

now i we combine like terms, i.e add \[A+\frac{1}{2}A +A\]

OpenStudy (anonymous):

our real job is to add \[1+\frac{1}{2}+1=2+\frac{1}{2}=\frac{5}{2}\]

OpenStudy (anonymous):

that is where the \[\frac{5}{2}A\] came from, and why we write \[\frac{5}{2}A+1<66\] clear?